a) Egerton Sacco owns a property worth Kshs. 1.5 Million. The Sacco can sell this property at Ksh. 1 million to finance a new venture which will generate a new income of Kshs. 700,000 per year before tax. The same property will cost the Sacco Kshs. 200,000 per year to lease. The venture is a trading business whose life span is estimated at 15 years and corporate tax is 35%. Advice the Sacco which is the best alternative. (6 marks)

b) Explain the factors that you would consider when choosing between buying and leasing of property alternatives.


c) An investor purchased 50 hectares of land 10 years ago at Kshs 20,000 per hectare. He incurred and expense of Kshs, 15,000 per hectare on clearing and planting tea on the land. He has been spending Kshs. 1,000 per year to look after the tea farm. Assuming a rate of interest of 8%, at what price should he sell this farm for without making loss? (5 marks)


d) Explain three reasons why property buyers and sellers may prefer to use brokerage services rather than to go it alone.
Answer #1

a) If Egerton Sacco opts to sell this property @ 1million to finance the new venture earning 7,00,000 each year. It will also result in savings in cost of leasing the property.
Cash inflow from sale @ Year 0 = Ksh 10,00,000
Net Cash inflow from Year 1-15 = (7,00,000 - 2,00,00)(1-0.35) * 15 = 3,25,000*15 = 48,75,000
Therefore, the total Cash inflow accruing due to such project is Kshs 58,75,000 over the 15 years time. However the actual figure should be discounted. In absence of any such figure figures are compared without discounting.

Since cashflows from option of selling to finance new venture is more than the value of property, Such option of selling to finance the new venture should be opted by Egerto Sacco.

b) The factors to be considered when choosing between buying or leasing options are as follows:
1. Location of the property - Considering the long term view: For certain types of businesses(trading), location would be considered to be an important factor as the quantum of sales will be directly affected by the location of the property. Buying property at an appropriate location will ensure that the premises need not be shifted due to non renewal of leases. Buying option would be preferable in such scenarios.
2. Cash flow : Cash flow requirement in case of leases is significantly lower compared to buying option as leasing only requires security deposit and monthly/quarterly/semi annual/annual lease payments.
3. Tax Savings: Tax Savings due to lease payments & depreciation to be taken into consideration.
4: Net Present Value: The net present value of all the payments made under leasing option to be compared with the buying decision to decide whether to lease or buy the property.

c)
Current value of  Expenditure incurred @ year 0(ie 10 year ago) compounded @ 8% for 10 years
= (20,000* 50) + (15000*50) (1+0.08)^10 = Kshs. 37,78,118.75

Current value of expenditure incurred every year for 10 years( Assuming expenditure is incurred at the start of the year)
= 1000*(15.64544874629) = Kshs. 15645.45
Therefore he should sell this farm at least at 37,78,118.75 + 15645.45 = Kshs. 37,93,764.2